1. solve for n: 32n-256=-144
2. solve for x: -3x+198=-303
3. If each school bus carries most 20 students, and the school used 5 busess ;last Monday, then which statements shows how many (n) students cam to school by bus?
a. n~<100 b. n=100 c.N>100 d. n<100
4. Don checked the thermometer in his art studio. It read 68 degrees.
What is the temperature in degrees celsius?

Answer by Nate(3500)

1. solve for n: 32n-256=-144
32n-256=-144 original
32n=112 add 256 to each side of the equal sign
n=3.5 divide 32 throughout the equation
-------------------------------------------------------------------------------
2. solve for x: -3x+198=-303
-3x+198=-303 original
-3x=-501 subtract 198 from each side
x=167 divide by -3
-------------------------------------------------------------------------------
3. If each school bus carries most 20 students, and the school used 5 busess ;last Monday, then which statements shows how many (n) students cam to school by bus?
a. n~<100 b. n=100 c.N>100 d. n<100
We know that the max students to ride the buses would be 100 students so c and d are not right! Also, some students can be sick, so all 100 students will not ride the bus every day so b isn't right. The answer is a!
-------------------------------------------------------------------------------
4. Don checked the thermometer in his art studio. It read 68 degrees.
What is the temperature in degrees celsius?
the equation is: c=(5/9)(F-32)
c=(5/9)(F-32)
c=(5/9)(68-32)
c=(5/9)(36)
c=20
20 degrees celsius!